# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

### v1

@@ 2.5 Efficiency of the estimator $\hat{\beta}$
As detailed out in the introduction, the estimator (9) coincides with the semiparametrically efficient estimator in the homoscedastic case (see for example equation (9) in [31]). For completeness, we provide a direct argument by comparing the asymptotic variance of $\hat{\beta}$ with the asymptotic variance of an estimator based on the identification (5). Such an estimator $\beta_{\zeta}$ is obtained by replacing $R_{f,i}^{k}$ in (9) by $R_{\zeta,i}^{k}$, where $R_{\zeta,i}^{k}=\hat{\zeta}^{k}(Z_{i},X_{i})-\hat{\mu}^{k}(X_{i})$, with $\hat{\mu}^{k}(X_{i})$ being an estimate of $\mu(X_{i})=\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$ and $\hat{\zeta}^{k}$ is either estimated or known a priori, as for example in the standard DML with linear instruments [12, 16] which is recovered by setting $\zeta(Z_{i},X_{i})=Z_{i}$. Analogously to $\hat{\beta}$, one can show that $\hat{\beta}_{\zeta}$ is asymptotically normal with asymptotic variance given by
@@ 3 Heterogeneous Treatment Effect
$Y_{i}=\beta(A_{i})D_{i}+g(X_{i})+\epsilon_{i}.$ (21)
The only change compared to the model (3) is that the treatment effect is assumed to be heterogeneous with respect to the univariate continuous random variable $A_{i}$ and we again assume the availability of an instrumental variable $Z_{i}\in\mathbb{R}^{d}$. As for the homogeneous treatment effect, we allow the data generating process to vary with $N$, even if it is not reflected in the notation.
The goal is to perform inference for $\beta(a)$, that is, the value of the treatment effect for a particular value of $A_{i}=a\in\mathbb{R}$. We need the following extension of Assumption 1 to make $\beta(a)$ identifiable.
@@
    $\mathbb{E}\left[\left(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}]\right)^{2}|A_{i}=a\right]\neq 0$,
@@
    $A_{i}$ is measurable with respect to the sigma-algebra generated by $X_{i}$,
@@
    $A_{i}$ is absolutely continuous with respect to the Lebesgue measure and has density $p_{A}(\cdot)$.
Assertion 1 encodes that $Z_{i}$ is a relevant IV given $X_{i}$ conditional on $A_{i}=a$. Assertion 2 is the same as for Assumption 1 and states that $Z_{i}$ is a valid instrumental variable. Assertion 3 states that $A_{i}$ is measurable with respect to $X_{i}$. Usually, $A_{i}$ is just a coordinate of $X_{i}$ and in fact, assertion 3 of Assumption 2 can be made without loss of generality since we can otherwise just replace $X_{i}\leftarrow(X_{i},A_{i})$. Under Assumption 2, we can identify $\beta(a)$ by simply making the identification from the homogeneous case conditional on $A_{i}=a$, i.e.
$\beta(a)=\frac{\mathbb{E}\left[(Y_{i}-\mathbb{E}[Y_{i}|X_{i}])(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])|A_{i}=a\right]}{\mathbb{E}\left[(D_{i}-\mathbb{E}[D_{i}|X_{i}])(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])|A_{i}=a\right]}.$ (22)
@@ 3.1 Estimators
$\hat{\beta}(a)=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{f,i}^{k}K\left(\frac{A_{i}-a}{h}\right)}{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{A_{i}-a}{h}\right)}.$ (23)
The main intuition for (23) is that we can view the treatment effect as being homogeneous when we restrict the attention to $A_{i}$ being near the target level $a$. In a similar fashion to the homogeneous treatment effect, we will show below in Section 3.3 that $\sqrt{Nh}(\hat{\beta}(a)-\hat{\beta}(a))/\sigma(a)\to\mathcal{N}(0,1)$ with asymptotic variance $\sigma^{2}(a)$ that can be consistently estimated by
$\hat{\sigma}^{2}(a)=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\hat{\beta}(a)R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}}{\left(\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right)^{2}}.$ (24)
@@
As for the homogeneous treatment effect, we recommend using a two-stage approach for the residuals $R_{f,i}^{k}$, see Remark 2 for the details.
Remark 6.
As for the homogeneous treatment effect, in principle one can replace $R_{f,i}^{k}$ by $R_{\zeta,i}^{k}=\hat{\zeta}^{k}(Z_{i},X_{i})-\hat{\mu}^{k}(X_{i})$ for an arbitrary function $\zeta(Z_{i},X_{i})$ and $\hat{\mu}^{k}(X_{i})$ being an estimate of $\mu(X_{i})=\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$. Most prominently, one could take the “linear IV” estimator $\zeta(Z_{i},X_{i})=Z_{i}$ and $\mu(X_{i})=\mathbb{E}[Z_{i}|X_{i}]$. However, when using the same kernel and bandwidth, the estimator based on $\zeta(Z_{i},X_{i})=f(Z_{i},X_{i})$ is the most efficient one under homoscedasticity, as we will see in Remark 11.
@@ 3.2 Robust Confidence Sets
$\hat{Q}(\gamma,a)=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\gamma R_{D,i}^{k}\right)R_{f,i}^{k}K\left(\frac{A_{i}-a}{h}\right),$
@@
$\widehat{SE}^{2}_{Q}(\gamma,a)=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\gamma R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}-h\hat{Q}(\gamma,a)^{2}.$
Below in Section 3.3.1, we will give conditions such that under $H_{0}:\beta(a)=\beta_{0}(a)$,
$\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(a),a)}{\widehat{SE}_{Q}(\beta_{0}(a),a)}\to\mathcal{N}(0,1),$
and consequently, $\mathcal{C}_{N}^{\alpha}(a)\coloneqq\left\{\gamma:|\hat{Q}(\gamma,a)|\leq z_{1-\alpha/2}\widehat{SE}_{Q}(\beta)/\sqrt{Nh}\right\}$ is an asymptotic $(1-\alpha)$-confidence set for $\beta(a)$, similarly to the homogeneous case.
@@
In Appendix C.1, we derive an explicit expression for the robust confidence sets and in Appendix C.2, we show that under certain conditions on the model, the robust confidence sets are asymptotically equivalent to standard confidence intervals based on $\hat{\sigma}^{2}(a)$. This means that one does not lose asymptotic efficiency by using the robust confidence sets.
@@ 3.3 Theory
In classical univariate kernel smoothing, the optimal bandwidth satisfies $h\asymp N^{-1/5}$. This yields the smallest mean squared error of the nonparametric estimator, since with this choice, the bias and the variance are of the same order. However, to obtain valid inference for $\beta(a)$, we need that the variance dominates the bias and hence, one needs to choose a smaller bandwidth (“undersmoothing”) [34]. This is also true for our setting and we need the following assumption.
@@
The remaining assumptions can be divided into conditions on the convergence rates of the nuisance function estimates and some more technical (mainly smoothness) assumptions needed to apply the kernel smoothing successfully. We start with the assumptions on the nuisance function estimates. In contrast to the usual conditions for double/debiased machine learning that essentially require that all the nuisance functions converge faster than $N^{-1/4}$ in $L_{2}$-norm (see condition (16) above), we need slightly faster convergence requirements on the nuisance functions, which we outline below. The reason for this is that we only want to make assumptions on the overall convergence rate of the nuisance functions, e.g., $\|\hat{f}-f\|_{L_{2}}$ and not make any assumptions on convergence rates that are somehow localized/conditional on $A_{i}=a$.
@@
$\sigma_{Q}(a)^{2}=\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a]p_{A}(a)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x,$ (25)
where we recall that $p_{A}(\cdot)$ is the density of $A_{i}$ with respect to the Lebesgue measure. Similarly to the homogeneous treatment effect, $\sigma_{Q}(a)^{2}$ can be seen as an unnormalized measure of IV strength conditional on $A_{i}=a$ on the population level.
@@
    For $k=1,\ldots,K$, it holds that $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}(a)=o_{P}\left(N^{-1/2}h^{-1}\right)$ and $\|\phi-\hat{\phi}^{k}\|_{L_{2}}/\sigma_{Q}(a)=o_{P}\left(N^{-1/2}h^{-1}\right)$.
@@
    For $k=1,\ldots,K$, it holds that $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}(a)=o_{P}\left(h^{1/2}\right)$ and $\|\phi-\hat{\phi}^{k}\|_{L_{2}}/\sigma_{Q}(a)=o_{P}\left(h^{1/2}\right)$.
@@
    $\frac{1}{\sigma_{Q}(a)}\|f-\hat{f}^{k}\|_{L_{2}}\|l-\hat{l}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(a)}\|\phi-\hat{\phi}^{k}\|_{L_{2}}\|l-\hat{l}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(a)}\|f-\hat{f}^{k}\|_{L_{2}}\|\phi-\hat{\phi}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(a)}\|\phi-\hat{\phi}^{k}\|_{L_{2}}^{2}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right).$
@@
    There exist $p_{1},q_{1}\in[1,\infty]$ with $11/p_{1}+1/q_{1}=$ 1 and $C<\infty$ independent of $N$ such that $\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p_{1}}}/\sigma_{Q}^{2}(a)\leq C$ and for $k=1,\ldots,K$ it holds that $\|l-\hat{l}^{k}\|_{L_{2q_{1}}}=o_{P}\left(h^{1/2}\right)$.
@@
    There exist $p_{2},q_{2}\in[1,\infty]$ with $11/p_{2}+1/q_{2}=$ 1 and $C<\infty$ independent of $N$ such that $\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p_{2}}}/\sigma_{Q}^{2}(a)\leq C$ and for $k=1,\ldots,K$, it holds that $\|\phi-\hat{\phi}^{k}\|_{L_{2q_{2}}}=o_{P}\left(h^{1/2}\right)$.
@@
Assume that $\sigma_{Q}(a)^{2}\asymp 1$, which is for example the case if the data generating process (21) does not change with $N$. From Assumption 4, we know that $h\ll N^{-1/5}$. Hence, assertion 1 of Assumption 5 is satisfied if $\|f-\hat{f}^{k}\|_{L_{2}}=O_{P}(N^{-3/10})$ and $\|\phi-\hat{\phi}^{k}\|_{L_{2}}=O_{P}(N^{-3/10})$, which is a slightly stronger requirement than $o_{P}\left(N^{-1/4}\right)$. The second assertion is comparably weak. The third assertion are the conditions on the products of the $L_{2}$ errors of the nuisance function estimators. Since $h\ll N^{-1/5}$, these assumptions are stronger than assuming that the respective products are $o_{P}\left(N^{-3/5}\right)$. Hence, the assumptions on the cross terms are also slightly stronger than for the homogeneous treatment effect.
When $\sigma_{Q}^{2}(a)\to 0$, this corresponds to models with weak IV. As already discussed in Remark 3 we cannot allow $\sigma_{Q}^{2}(a)$ to go to zero too fast since otherwise, the requirements on the convergence rates for the nuisance functions cannot be satisfied.
@@
    $p_{A}(a)$ is bounded away from 00.
@@
    If $\sigma_{Q}^{2}(a)\to 0$ as $N\to\infty$, this is attributed to weak IV, i.e. $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a]\to 0$ and not to other changes in the data generating process (for example not because $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\to 0$).
@@
    Various functions $x\mapsto\rho(x)$ are twice differentiable for $x$ in the support of $A_{i}$ with bounded second derivative (when properly rescaled by $\sigma_{Q}(a)$). The functions $\rho(\cdot)$ are products of $\beta(\cdot)$, $p_{A}(\cdot)$ and various conditional moments as for example $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=\cdot]$.
With these assumptions, we obtain the asymptotic normality of the estimator $\hat{\beta}(a)$.
@@
Under Assumptions 2, 3, 4, 5 and 6, additionally assume that $\sigma_{Q}^{2}(a)\gg(Nh)^{-1}$, $\frac{\|f-\hat{f}^{k}\|_{L_{2}}}{\sigma_{Q}^{2}(a)}=o_{P}\left(h^{1/2}\right)$ and $\frac{\|\phi-\hat{\phi}^{k}\|_{L_{2}}}{\sigma_{Q}^{2}(a)}=o_{P}\left(h^{1/2}\right)$. Then, we have
$\frac{\sqrt{Nh}(\hat{\beta}(a)-\beta(a))}{\sigma(a)}\to\mathcal{N}(0,1)\text{ in distribution,}$,
@@
$\sigma^{2}(a)=\frac{\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a]\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x}{\mathbb{E}[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))|A_{i}=a]^{2}p_{A}(a)}.$
The following theorem ensures that the estimator $\hat{\sigma}^{2}(a)$ defined in (24) is a consistent estimator of $\sigma^{2}(a)$ hence one can construct asymptotic confidence intervals for $\beta(a)$.
@@
Additionally to the conditions of Theorem 9, assume that $\frac{\|f-\hat{f}^{k}\|_{L_{4}}}{\sigma_{Q}(a)}=O_{p}\left(h^{1/4}\right)$ and $\frac{\|\phi-\hat{\phi}^{k}\|_{L_{4}}}{\sigma_{Q}(a)}=O_{p}\left(h^{1/4}\right)$. Then, we have
$\frac{\hat{\sigma}^{2}(a)}{\sigma^{2}(a)}\to 1\text{ in probability.}$.
@@
Regarding the efficiency of the estimator $\hat{\beta}(a)$, we can exactly follow the argument from the proof of Proposition 4. Concretely, given a fixed kernel $K$, fixed bandwidth $h$, there is no transformation $\zeta(Z_{i},X_{i})$ such that the resulting estimator according to Remark 6 is more efficient than our estimator $\hat{\beta}(a)$ under homosecadasticity $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathrm{const.}$

### v2

@@ 2.5 Efficiency of the estimator $\hat{\beta}$
As detailed out in the introduction, the estimator (10) coincides with the semiparametrically efficient estimator in the homoscedastic case (see for example equation (9) in [47]). For illustration, we directly compare the asymptotic variance of $\hat{\beta}$ with the asymptotic variance of an estimator based on the identification (6). Such an estimator $\beta_{\zeta}$ is obtained by replacing $R_{f,i}^{k}$ in (10) by $R_{\zeta,i}^{k}$, where $R_{\zeta,i}^{k}=\hat{\zeta}^{k}(Z_{i},X_{i})-\hat{\mu}^{k}(X_{i})$, with $\hat{\mu}^{k}(X_{i})$ being an estimate of $\mu(X_{i})=\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$ and $\hat{\zeta}^{k}$ is either estimated or known a priori, as for example in the standard DML with linear instruments [16, 21] which is recovered by setting $\zeta(Z_{i},X_{i})=Z_{i}$. Analogously to $\hat{\beta}$, one can show that $\hat{\beta}_{\zeta}$ is asymptotically normal with asymptotic variance given by
@@ 3 Heterogeneous Treatment Effect
$Y_{i}=\beta(V_{i})D_{i}+g(X_{i})+\epsilon_{i}.$ (21)
The only change compared to the model (4) is that the treatment effect is assumed to be heterogeneous with respect to the univariate continuous random variable $V_{i}$ and we again assume the availability of an instrumental variable $Z_{i}\in\mathbb{R}^{d}$. As for the homogeneous treatment effect, we allow the data generating process to vary with $N$, even if it is not reflected in the notation.
The goal is to perform inference for $\beta(v)$, that is, the value of the treatment effect for a particular value of $V_{i}=v\in\mathbb{R}$. We need the following extension of Assumption 1 to make $\beta(v)$ identifiable.
@@
    $\mathbb{E}\left[\left(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}]\right)^{2}|V_{i}=v\right]\neq 0$,
@@
    $V_{i}$ is measurable with respect to the sigma-algebra generated by $X_{i}$,
@@
    The distribution of $V_{i}$ is absolutely continuous with respect to the Lebesgue measure and has density $p_{V}(\cdot)$.
Assertion 1 encodes that $Z_{i}$ is a relevant IV given $X_{i}$ conditional on $V_{i}=v$. Assertion 2 is the same as for Assumption 1 and states that $Z_{i}$ is a valid instrumental variable. Assertion 3 states that $V_{i}$ is measurable with respect to $X_{i}$. Usually, $V_{i}$ is just a coordinate of $X_{i}$ and in fact, assertion 3 of Assumption 2 can be made without loss of generality since we can otherwise just replace $X_{i}\leftarrow(X_{i},V_{i})$. The absolute continuity of the distribution of $V_{i}$ with respect to the Lebesgue measure in assertion 4 is assumed for the sake of conciseness of the exposition. An analogous approach would work for discrete random variables $V_{i}$, which would reduce to simple binning. Under Assumption 2, we can identify $\beta(v)$ by simply making the identification from the homogeneous case conditional on $V_{i}=v$, i.e.
$\beta(v)=\frac{\mathbb{E}\left[(Y_{i}-\mathbb{E}[Y_{i}|X_{i}])(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])|V_{i}=v\right]}{\mathbb{E}\left[(D_{i}-\mathbb{E}[D_{i}|X_{i}])(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])|V_{i}=v\right]}.$ (22)
@@ 3.1 Estimators
$\hat{\beta}(v)=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)}{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)}.$ (23)
The main intuition for (23) is that we can view the treatment effect as being homogeneous when we restrict the attention to $V_{i}$ being near the target level $v$. In a similar fashion to the homogeneous treatment effect, we will show below in Section 3.3 that $\sqrt{Nh}(\hat{\beta}(v)-\hat{\beta}(v))/\sigma(v)\to\mathcal{N}(0,1)$ with asymptotic variance $\sigma^{2}(v)$ that can be consistently estimated by
$\hat{\sigma}^{2}(v)=\frac{\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\hat{\beta}(v)R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}}{\left(\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right)^{2}}.$ (24)
@@
As for the homogeneous treatment effect, in principle one can replace $R_{f,i}^{k}$ by $R_{\zeta,i}^{k}=\hat{\zeta}^{k}(Z_{i},X_{i})-\hat{\mu}^{k}(X_{i})$ for an arbitrary function $\zeta(Z_{i},X_{i})$ and $\hat{\mu}^{k}(X_{i})$ being an estimate of $\mu(X_{i})=\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$. Most prominently, one could take the “linear IV” estimator $\zeta(Z_{i},X_{i})=Z_{i}$ and $\mu(X_{i})=\mathbb{E}[Z_{i}|X_{i}]$. However, when using the same kernel and bandwidth, the choice $\zeta(Z_{i},X_{i})=f(Z_{i},X_{i})$ leads to the most efficient estimator of this form under homoscedasticity, as we will see in Remark 10.
@@ 3.2 Robust Confidence Sets
$\hat{Q}(\gamma,v)=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\gamma R_{D,i}^{k}\right)R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right),$
@@
$\widehat{SE}^{2}_{Q}(\gamma,v)=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}-\gamma R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}-h\hat{Q}(\gamma,v)^{2}.$
Below in Section 3.3.1, we will give conditions such that under $H_{0}:\beta(v)=\beta_{0}(v)$,
$\frac{\sqrt{Nh}\hat{Q}(\beta_{0}(v),v)}{\widehat{SE}_{Q}(\beta_{0}(v),v)}\to\mathcal{N}(0,1),$
and consequently, $\mathcal{C}_{N}^{\alpha}(v)\coloneqq\left\{\gamma:|\hat{Q}(\gamma,v)|\leq z_{1-\alpha/2}\widehat{SE}_{Q}(\beta)/\sqrt{Nh}\right\}$ is an asymptotic $(1-\alpha)$-confidence set for $\beta(v)$, similarly to the homogeneous case.
@@
In Appendix C.1, we derive an explicit expression for the robust confidence sets and in Appendix C.2, we show that under certain conditions on the model, the robust confidence sets are asymptotically equivalent to standard confidence intervals based on $\hat{\sigma}^{2}(v)$. This means that one does not lose asymptotic efficiency by using the robust confidence sets.
@@ 3.3 Theory
In classical univariate kernel smoothing, the optimal bandwidth satisfies $h\asymp N^{-1/5}$. This yields the smallest mean squared error of the nonparametric estimator, since with this choice, the bias and the variance are of the same order. However, to obtain valid inference for $\beta(v)$, we need that the variance dominates the bias and hence, one needs to choose a smaller bandwidth (“undersmoothing”) [51]. This is also true for our setting and we need the following assumption.
@@
The remaining assumptions can be divided into conditions on the convergence rates of the nuisance function estimates and some more technical (mainly smoothness) assumptions needed to apply the kernel smoothing successfully. We start with the assumptions on the nuisance function estimates. In contrast to the usual conditions for double/debiased machine learning that essentially require that all the nuisance functions converge faster than $N^{-1/4}$ in $L_{2}$-norm (see condition (16) above), we need slightly faster convergence requirements on the nuisance functions, which we outline below. The reason for this is that we only want to make assumptions on the overall convergence rate of the nuisance functions, e.g., $\|\hat{f}-f\|_{L_{2}}$ and not make any assumptions on convergence rates that are somehow localized/conditional on $V_{i}=v$.
@@
$\sigma_{Q}^{2}(v)=\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]p_{V}(v)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x,$ (25)
where we recall that $p_{V}(\cdot)$ is the density of $V_{i}$ with respect to the Lebesgue measure. Similarly to the homogeneous treatment effect, $\sigma_{Q}^{2}(v)$ can be seen as an unnormalized measure of IV strength conditional on $V_{i}=v$ on the population level.
@@
    For $k=1,\ldots,K$, it holds that $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}(v)=o_{P}\left(N^{-1/2}h^{-1}\right)$ and $\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}/\sigma_{Q}(v)=o_{P}\left(N^{-1/2}h^{-1}\right)$.
@@
    For $k=1,\ldots,K$, it holds that $\|f-\hat{f}^{k}\|_{L_{2}}/\sigma_{Q}(v)=o_{P}\left(h^{1/2}\right)$ and $\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}/\sigma_{Q}(v)=o_{P}\left(h^{1/2}\right)$.
@@
    $\frac{1}{\sigma_{Q}(v)}\|f-\hat{f}^{k}\|_{L_{2}}\|l-\hat{l}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(v)}\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}\|l-\hat{l}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(v)}\|f-\hat{f}^{k}\|_{L_{2}}\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right),$
    $\frac{1}{\sigma_{Q}(v)}\|\phi-\hat{\phi}^{k}_{1}\|_{L_{2}}\|\phi-\hat{\phi}^{k}_{2}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}h^{1/2}\right).$
@@
    There exist $p_{1},q_{1}\in[1,\infty]$ with $1/p_{1}+1/q_{1}=1$ and $C<\infty$ independent of $N$ such that $\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p_{1}}}/\sigma_{Q}^{2}(v)\leq C$ and for $k=1,\ldots,K$ it holds that $\|l-\hat{l}^{k}\|_{L_{2q_{1}}}=o_{P}\left(h^{1/2}\right)$.
@@
    There exist $p_{2},q_{2}\in[1,\infty]$ with $1/p_{2}+1/q_{2}=1$ and $C<\infty$ independent of $N$ such that $\|\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|X_{i}]\|_{L_{p_{2}}}/\sigma_{Q}^{2}(v)\leq C$ and for $k=1,\ldots,K$, it holds that $\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2q_{2}}}=o_{P}\left(h^{1/2}\right)$.
@@
Assume that $\sigma_{Q}^{2}(v)\asymp 1$, which is for example the case if the data generating process (21) does not change with $N$. From Assumption 4, we know that $h\ll N^{-1/5}$. Hence, assertion 1 of Assumption 5 is satisfied if $\|f-\hat{f}^{k}\|_{L_{2}}=O_{P}(N^{-3/10})$ and $\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}=O_{P}(N^{-3/10})$, which is a slightly stronger requirement than $o_{P}\left(N^{-1/4}\right)$. The second assertion is comparably weak. The third assertion are the conditions on the products of the $L_{2}$ errors of the nuisance function estimators. Since $h\ll N^{-1/5}$, these assumptions are stronger than assuming that the respective products are $o_{P}\left(N^{-3/5}\right)$. Hence, the assumptions on the cross terms are also slightly stronger than for the homogeneous treatment effect.
When $\sigma_{Q}^{2}(v)\to 0$, this corresponds to models with weak IV. As already discussed in Remark 3 we cannot allow $\sigma_{Q}^{2}(v)$ to go to zero too fast since otherwise, the requirements on the convergence rates for the nuisance functions cannot be satisfied.
@@
    $p_{V}(v)$ is bounded away from 0.
@@
    If $\sigma_{Q}^{2}(v)\to 0$ as $N\to\infty$, this is attributed to weak IV, i.e. $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]\to 0$ and not to other changes in the data generating process (for example not because $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]\to 0$).
@@
    Various functions $x\mapsto\rho(x)$ are twice differentiable for $x$ in the support of $V_{i}$ with bounded second derivative (when properly rescaled by $\sigma_{Q}(v)$). The functions $\rho(\cdot)$ are products of $\beta(\cdot)$, $p_{V}(\cdot)$ and various conditional moments as for example $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=\cdot]$.
With these assumptions, we obtain the asymptotic normality of the estimator $\hat{\beta}(v)$.
@@
Under Assumptions 2, 3, 4, 5 and 6, additionally assume that $\sigma_{Q}^{2}(v)\gg(Nh)^{-1}$, $\frac{\|f-\hat{f}^{k}\|_{L_{2}}}{\sigma_{Q}^{2}(v)}=o_{P}\left(h^{1/2}\right)$ and $\frac{\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}}{\sigma_{Q}^{2}(v)}=o_{P}\left(h^{1/2}\right)$. Then, we have
$\frac{\sqrt{Nh}(\hat{\beta}(v)-\beta(v))}{\sigma(v)}\to\mathcal{N}(0,1)\text{ in distribution,}$
@@
$\sigma^{2}(v)=\frac{\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x}{\mathbb{E}[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))|V_{i}=v]^{2}p_{V}(v)}.$
The following theorem ensures that the estimator $\hat{\sigma}^{2}(v)$ defined in (24) is a consistent estimator of $\sigma^{2}(v)$ hence one can construct asymptotic confidence intervals for $\beta(v)$.
@@
Additionally to the conditions of Theorem 8, assume that $\frac{\|f-\hat{f}^{k}\|_{L_{4}}}{\sigma_{Q}(v)}=O_{p}\left(h^{1/4}\right)$ and $\frac{\|\phi-\hat{\phi}_{2}^{k}\|_{L_{4}}}{\sigma_{Q}(v)}=O_{p}\left(h^{1/4}\right)$ Then, we have
$\frac{\hat{\sigma}^{2}(v)}{\sigma^{2}(v)}\to 1\text{ in probability.}$
@@
Regarding the efficiency of the estimator $\hat{\beta}(v)$, we can exactly follow the argument from the proof of Proposition 4. Concretely, given a fixed kernel $K$, fixed bandwidth $h$, there is no transformation $\zeta(Z_{i},X_{i})$ such that the resulting estimator according to Remark 5 is more efficient than our estimator $\hat{\beta}(v)$ under homoscedasticity $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathrm{const.}$